# KIX: A Metacognitive Generalization Framework

### v1

@@ Abstract
Humans and other animals aptly exhibit general intelligence behaviors in solving a variety of tasks with flexibility and ability to adapt to novel situations by reusing and applying high level knowledge acquired over time. But artificial agents are more of a specialist, lacking such generalist behaviors. Artificial agents will require understanding and exploiting critical structured knowledge representations. We present a metacognitive generalization framework, Knowledge-Interaction-eXecution (KIX), and argue that interactions with objects leveraging type space facilitate the learning of transferable interaction concepts and generalization. It is a natural way of integrating knowledge into reinforcement learning and promising to act as an enabler for autonomous and generalist behaviors in artificial intelligence systems.
@@ 1 Introduction
Figure 1: Learning for generalist behaviors necessitates building a natural yet abstract understanding of the environment, as well as meta-reasoning and interaction abilities. The metacognitive knowledge-guided approach captures the inherent abstractions and relations between objects to generate interaction goals, enabling the agent to interact with objects and generalize to solve complex tasks.
@@
Flexibility, ability to actively think and evaluate options –learning and reasoning to enable adaptation– are hallmarks of human intelligence. Core to this intelligence is how easily people can find natural decomposition in world environments and tasks, enabling flexible, generalist behaviors. Replicating such autonomy and general intelligence abilities in artificial agents continues to pose significant challenges, which we argue requires a knowledge-centric metacognitive approach.
@@
Meta-reasoning about next actions, action values, and possible next states facilitates flexible and transferable decision making for novel situations, problem solving, concept learning, and knowledge acquisition. This reasoning involves more than just memory lookup. Meta-reasoning is more broadly sense-making in a knowledge system – enabling people to intuitively direct their learning to make sense of relationships between entities, interact with the same entities in different ways, and ground their interactions for both short and long-term goals. Metacognitive reasoning appears to be central to generalization. In contrast, meta-reasoning systems in artificial intelligence are more focused on resource allocation problems. Although resource allocation underlies efficient functioning, we need to ideate metacognitive generalization, figure 1, to enable generalist behaviors in artificial agents. Generalist behaviors have key characteristics which both constitute a basis of devising a metacognitive framework and compartmentalize the behaviors in a useful way.
@@
Our proposal builds on three main principles for solving complex tasks in a flexible manner. Natural decomposition and abstraction: World environments are naturally composed of entities and categorized by their types. It is similar to how a natural hierarchy of objects and classes is formed. Agents interact with entities in the instance space, which inherit properties and relationships from a higher-level type space. Interactions: World environments are essentially knowledge graphs that represent the underlying objects and their relationships. In order to complete its tasks and achieve its goals, an agent’s metacognitive process employs these knowledge graphs to create a plan that specifies which entities it must interact with and what interactions it should have. Composition of interactions: Agents’ flexible behaviors stem from their ability to learn interactions and leverage interaction pathways for graph traversal at the type-space and instance levels as needed for quick adaptation. Integrating these principles into metacognitive learning, reasoning, and control provides levers and connecting points for instilling generalist behaviors in agents.
@@
Our proposed meta-cognitive framework captures the essence of generalist behaviors and may have far-reaching implications across domains. In this article, we provide a description of the key components of the proposed framework, which exemplify generalist characteristics followed by experimental evaluation, as well as how we envision it enabling flexibility through meta-level knowledge guided interactions.
@@ 2 Background
In this section, we briefly describe closely related work as well as the general setting of meta-reasoning, graph neural networks, and policy gradient methods.
@@
Related work Compositionality and abstractions are central for generalization learning. Abstraction requires a method for grounding representations to the actual objects. In previous work, Veerapaneni et al. (2020) presented an object-centric model in which object representations were constructed from raw observations and local simpler functions with uniformity, and then used for prediction and planning. Similarly, Vijay et al. (2019) used a symbolic object-centric representation to study the effect of adding prior knowledge through relations in an object graph. These works only consider object space, which inhibits the abstraction of objects and relational structures that occur naturally in the real world and are used for making decisions. Some of the main questions are what an abstraction should entail and how it can be used to compose interactions that can be used as building blocks for leveraging natural relational structures. We attempt to address these questions, which have been largely overlooked in previous works. Moreover, what does natural abstraction mean? It implies preserving relational higher-order structures and abstractions that occur naturally in the real world. Our proposal preserves abstract knowledge in the form of knowledge graphs, where abstractions are grounded using type space. Learning a compositional structure is accomplished by learning over knowledge graphs, with neighborhoods aggregated until stability is reached. Different from previous works, that’s what our proposal entails.
@@
Other methods emphasize aspects which are different from ours, such as encouraging exploration through intrinsic reward or information gain Ten et al. (2022); Oudeyer et al. (2007); Pathak et al. (2017), or motivation Moerland et al. (2023); Zhu et al. (2023); exploring for a variety of experiences by developing internal goals for empowerment Klyubin et al. (2005) ; using a curriculum to progress from easier to more difficult tasks Soviany et al. (2022); Campero et al. (2020) or inducing a learning hierarchy either based on teacher signals Portelas et al. (2020) Pateria et al. (2021); Hutsebaut-Buysse et al. (2022) or rewards Furelos-Blanco et al. (2023). These methods focus on the execution space, influencing reward signals for learning policies in one way or another but largely ignore the higher order structures and relations that objects in the environment have. Our work distinguishes itself by incorporating natural relational structures in the form of knowledge graphs and allowing such meta-level knowledge structures to guide interactions that govern low-level execution.
@@
Meta-reasoning Meta-reasoning is often defined as the set of processes that monitor our thought processes and regulate resource allocation Ackerman and Thompson (2017). Its basic structure Nelson (1990) consists of two levels: meta and object. The meta level’s functions include control and monitoring. The object’s state changes as control passes from the meta to the object level. The monitoring process allows the meta controller to track the status of object-level processes and actions, as well as the consequences of those actions. However, meta-cognitive reasoning in its current form as a resource manager is insufficient to exhibit characteristics of generalist behaviors, necessitating generalist meta-cognition approaches such as the one proposed here.
@@
Graph neural networks A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is defined by a set of nodes $\mathcal{V}$ with node attributes $\mathtt{X}_{v}$ for $v\in\mathcal{V}$ and a set of edges $\mathcal{E}$ between nodes. Edge from node $u\in\mathcal{V}$ to node $v\in\mathcal{V}$ is given by $(u,v)\in\mathcal{E}$ with attributes $\mathtt{e}_{u,v}$. One typical task is to predict the label or value of the entire graph. Graph neural networks (GNN) (Gilmer et al. (2017); Hamilton et al. (2017) and others) use node connectivity as well as node and edge features to learn a permutation-invariant embedding vector for each node $h_{v},v\in\mathcal{V}$ and the entire graph $h_{\mathcal{G}}$. Typically, the representation is learned through message passing or neighborhood feature aggregation, in which node embedding vectors are updated in each layer based on message aggregation from neighbors:
@@
where $h_{v}^{(k)}$ is the updated representation of node $v$ embedding in layer $k$, $\mathcal{N}(v)$ is the set of neighboring nodes of node $v$, $h_{u}^{(k-1)}$ is the representation of neighboring node $u$ in layer $k-1$, $\bigoplus$ is a differentiable, permutation invariant aggregation function, $\mathcal{F}^{\prime}$ and $\mathcal{F}^{\prime\prime}$ are differentiable functions usually neural networks. Node features are transformed prior to aggregation, and the methods used to transform and aggregate features differ between graph network variants. Motivated by the attention mechanism Vaswani et al. (2017) in the transformer model, Veličković et al. (2017) employs self-attention over node features and aggregating over neighborhood nodes in graph attention networks (GAT). We use the GATv2 operator Brody et al. (2021), which appears to improve static attention problems in graph attention layers.
@@
Policy gradient methods We consider a standard reinforcement learning setting Sutton and Barto (2018) in which an agent interacts with the environment over a number of steps. A Markov decision process (MDP) is defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ in which $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite discrete action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the transition model, $r:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function and $\gamma\in(0,1)$ is the discount factor. At each step t, the agent takes an action $a_{t}\in\mathcal{A}$ in state $s_{t}\in\mathcal{S}$ according to its policy $\pi$, which is a mapping of states to actions. As a result, the agent reaches state $s_{t+1}\in\mathcal{S}$, receives a scalar reward $r_{t}\in\mathbb{R}$, and accumulates returns; $R_{t}=\sum_{k=0}^{\inf}\gamma^{k}r_{t+k+1}$ represents the accumulated return from step t. The agent’s goal is to maximize the expected return for each state. The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is defined as $Q^{\pi}(s,a)=\mathbb{E}[R_{t}|s_{t}=s,a]$, which is the expected return from selecting action $a$ in state $s_{t}$ following policy $\pi$. The value of state $s$ under policy $\pi$ is defined as $V_{\pi}(s)=\mathbb{E}[R_{t}|s_{t}=s]$, which is the expected return from the state $s$. Policy gradient methods, like REINFORCE Williams (1992) , directly learn a parameterized policy $\pi(s;\theta)$ and update its parameters $\theta$ using approximate gradient ascent of a performance measure, $\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi}[G_{t}\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})]$. To reduce the variance in the expected return estimate, a baseline is subtracted from it; typically, a learned value function is used as the baseline. The resulting quantity $A(s,a)=Q(s,a)-V(s)$ can be viewed as an estimate of the advantage of action $a$ in state $s$. This method is called advantage actor critic (A2C) Sutton and Barto (2018); Degris et al. (2012), and the agent acting according to the policy $\pi$ is the actor, while the value baseline is the critic. Several improved methods have been developed to address sample efficiency, stability, and training difficulty. Proximal policy optimization (PPO) Schulman et al. (2017) is an on-policy policy optimization technique that addresses concerns in actor critic methods using clipped objectives. We adapt PPO for learning policies.
@@ 3 Model
Our proposal for meta-cognitive generalization centers around the ideas of natural decomposition of tasks that an agent can accomplish through interactions with objects. Meta-policy recommends localized agent-object interactions. The agent executes the interactions following interaction policies and constructs a compositional solution. Here, we discuss the main ideas of the framework (see figure 2) and conjecture that it aptly exhibits task generalization characteristics.
@@
Figure 2: The meta-cognitive generalization framework consists of three levels: meta-level knowledge (K), interaction (I), and execution (X), which are based on levels of abstraction and computation to provide generalization capabilities. The knowledge or type space contains type space knowledge graphs representing higher order abstractions. The instance space contains object-level instance graphs and functions as an intermediary. The execution space is the lowest level abstraction in which the agent acts on meta-level recommendations. Computations proceed through recognition, recommendation, execution, and interactions to gather experiences for updates to interaction and meta policies.
Knowledge An agent can perform low-level actions $u=\{u_{1},u_{2},...\}$ in the environment $\Xi$. The environment consists of objects, which the agent’s perception systems associate with the entity types. Agent maintains its perceived knowledge in the form of knowledge graphs: a graph of objects and their relationships at the instance level, and a graph of entity types and their abstract relationships at the type level. A type space knowledge graph is an ordered pair $G_{K}=(E,\Lambda)$ that consists of a set of entity type nodes, $E=\{E_{1},E_{2},...,E_{m}\}$, and a set of edges or type level relations, $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$. The same type of relation between objects and abstract entity types has the same meaning; for example, adjacency relations indicate that two objects are adjacent to each other, whether in instance or type space. Using the knowledge graph and meta-level action of the meta-policy, $\lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{p}\}$, the meta-level of the agent determines how it will interact with an underlying object. To interact with an object, the agent activates an interaction type, $\lambda$, following the meta-policy guidance. For an interaction, the agent will perform a series of low-level actions. For example, if there is a door, the agent can activate ‘open the door’ at the abstract level, even though it requires many low-level actions such as moving and toggling. The agent uses an activation relation to keep track of which objects it wants to interact with and what type of interaction. Thus, the agent’s knowledge space $\mathcal{K}=(A,E,\Lambda,\lambda,G_{K})$ comprises of a type space knowledge graph, a set of entity types, abstract relations, and interactions.
Interaction The environment consists of objects, $e=\{e_{1},e_{2},...,e_{n}\}$, their types, $\text{type}(e_{i})\in E,\text{for}\ e_{i}\in e$, and is partially observable. Agent keeps track of the visible objects in its instance space and builds an instance graph, which is an ordered pair of nodes (observed objects) and edges (observed object relations), $G_{I}=(e,\Lambda_{I})$. Thus, the agent’s instance space, $\mathcal{I}=(A,e,\Lambda_{I},\lambda,G_{I})$, is made up of its instance space knowledge graph, relationships between objects, and available interactions. After constructing the knowledge graphs, the agent’s goal is to identify the objects for interaction which is done by the meta-level. Meta-level recommends an interaction goal for the agent. For the recommended interaction goals, the agent learns interaction policies for interacting with objects, as well as the relationships between interaction types and entity types. With the ability to interact with objects, the agent accumulates interaction knowledge to update the meta-level policy. Therefore, when viewing tasks as meta-policy guided interactions with objects, problems are naturally decomposed into interaction goals.
eXecution By pursuing interaction goals as directed by the meta-level, the agent learns to interact with objects and act at the execution level. The execution space for the agent is its partially observed view $X$. During an interaction, the agent takes actions $u$ on low-level states $s$. After learning both the interaction and meta policies, the agent acts in the environment, aided by interaction with objects. Therefore, unlike the models that are tightly coupled with low-level representations through end-to-end policies, agent actions are aided by meta-level knowledge-guided object interactions, providing a mechanism for generalist behaviors.
Mapping between execution space and instance space Assume that the agent has an object recognition system that can access its partially observed view $X$. The agent generates an instance-level graph using the observed objects. The function $\Phi:X\to G_{I}$ maps the agent’s object recognition to its instance graph by capturing observed objects and their relationships. As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view. The mapping function then generates the corresponding instance graph.
Mapping between instance space and type space The function $\Psi:G_{I}\to G_{K}$ provides a mapping between instance and type graphs. If the agent encounters a novel object for which it does not have a corresponding typed entity, it must either create a new type in its type space or assign the object to a type that is closely related. In the framework, $\Psi$ is learnable and represents a belief of mapping objects to their types. However, in this article, we treat the mapping as a fixed mapping from object nodes in $G_{I}$ to their corresponding type nodes in $G_{K}$.
@@
Proposal for activation object and interaction Using the meta-policy, the agent’s recommendation function $\Upsilon:G_{K},G_{I},\pi\to e_{i},\lambda_{p}$ recommends a candidate object $e_{i}$ for interaction and interaction goal $\lambda_{p}$. The recommender iterates through all visible objects in the agent’s instance graph $G_{I}$, then activates the object by adding an activation relationship between the agent and the object to the instance graph, as well as a corresponding activation relation between the agent and the object’s type node in the type space graph $G_{K}$. Activated $G_{K}$ represents the activated candidate object and serves as the input state for meta-policy, which returns both the state value and meta-action probabilities. The state value represents the value of activating the respective object. Thus, the recommender proposes an object for interaction $e_{i}$ by sampling from the values of activating objects, and the proposed interaction $\lambda_{p}$ is sampled from the meta-action probabilities of the proposed object. Acting on these recommendations, the agent collects interaction experiences $G_{K},\lambda$, that is, interaction trajectories $G_{K}^{0}\curvearrowright G_{K}^{1}...\curvearrowright G_{K}^{t}$, and learns meta-level policy.
Interactions At each step, the agent activates the object $e_{i}$ for interaction $\lambda_{p}$, which is its interaction goal, $g^{i}_{p}=(e_{i},\lambda_{p})$. Following this interaction goal, the agent interacts with the environment. After each round of interaction, the status of the interaction is passed to the meta level. Using the meta-policy, the meta-level recommender suggests next target object and interaction. The process continues until the task is completed or the environment is terminated. Throughout the process, the agent collects experience trajectories at both the interaction and meta levels. The policies are then updated based on the collected experiences.
@@
Meta-policy learning Meta-level representations of type space graphs with activation relations are learned using a message-passing graph neural network. Nodes have abstract level attributes such as type encoding, while edges have edge type encoding for different types of relationships. We use GAT convolutional layers Brody et al. (2021); Fey and Lenssen (2019) to learn graph representations given as,
@@
Interaction policy learning Interaction policies are a repository of policies for different interactions which can be recommended by the meta-level, $\mu_{\phi}=\{\mu_{\phi_{1}},\mu_{\phi_{2}},...,\mu_{\phi_{p}}\}$ and each policy is parameterized separately. When an interaction goal with an object $e_{i}$ and interaction $\lambda_{p}$ is recommended, $g^{i}_{p}=(e_{i},\lambda_{p})$, the agent attempts to achieve it by executing a series of low-level actions from $u$. The agent observes a low-level state $s$ from a convolutional neural network embedding, samples an action $u\sim\mu_{\phi_{p}}(.|s)$ using the respective interaction policy, acts in the environment, and collects the trajectory of experiences $\mathcal{D}_{\tau_{p}}$. The experiences collected are used to learn interaction policies. An estimate of advantages is computed as $A_{p}=r_{p}+\gamma_{p}V_{\mu_{\phi_{p}}}(s^{\prime})-V_{\mu_{\phi_{p}}}(s)$.

### v2

# KIX: A Knowledge and Interaction-Centric Metacognitive Framework for Task Generalization

@@ Abstract
People aptly exhibit general intelligence behaviors in solving a variety of tasks with flexibility and ability to adapt to novel situations by reusing and applying high-level knowledge acquired over time. But artificial agents are more like specialists, lacking such generalist behaviors. Artificial agents will require understanding and exploiting critical structured knowledge representations. We present a metacognitive generalization framework, Knowledge-Interaction-eXecution (KIX), and argue that interactions with objects leveraging type space facilitate the learning of transferable interaction concepts and generalization. It is a natural way of integrating knowledge into reinforcement learning and is promising to act as an enabler for autonomous and generalist behaviors in artificial intelligence systems.
@@ 1 Introduction
Figure 1: Learning generalist behaviors requires developing a natural and abstract understanding of the environment, as well as meta-reasoning and interaction abilities. The metacognitive knowledge-guided approach exploits inherent abstractions and relationships between objects to generate interaction goals, enabling the agent to interact with objects and generalize to solve complex tasks.
@@
Flexibility, ability to actively think and evaluate options –learning and reasoning to enable adaptation– are hallmarks of human intelligence. Core to this intelligence is how easily people can find natural decomposition in world environments and tasks; this enables flexible, generalist behaviors. Replicating such autonomy and general intelligence abilities in artificial agents continues to pose significant challenges, which we argue requires a knowledge and interaction-centric metacognitive approach.
@@
Meta-reasoning about next actions, their values, and possible next states facilitates flexible decision-making in novel situations, problem solving, concept learning, and knowledge acquisition. This reasoning extends beyond mere memory lookup. Meta-reasoning is more broadly sense-making in a knowledge system - enabling people to intuitively direct their learning to make sense of relationships between entities, interact with the same entities in diverse ways, and ground their interactions for both short- and long-term goals. Metacognitive reasoning appears to be central to generalization. In contrast, meta-reasoning systems in artificial intelligence are primarily focused on resource allocation problems. While resource allocation underpins efficient functioning, it is crucial to ideate metacognitive generalization, Figure 1, to enable generalist behaviors in artificial agents.
@@
Our proposal builds on three main principles for solving complex tasks flexibly. Natural Decomposition and Abstraction: World environments are naturally composed of entities categorized by their types, similar to a natural hierarchy of objects and classes. Agents interact with entities in the instance space, which inherit properties and relationships from a higher-level type space. Interactions: World environments are essentially knowledge graphs representing underlying objects and their relationships. An agent’s metacognitive process employs these knowledge graphs to create a plan specifying which entities to interact with and the nature of these interactions to complete tasks and achieve goals. Composition of Interactions: Agents’ flexible behaviors stem from their ability to learn interactions and leverage interaction pathways for graph traversal at both the type level and the instance level, enabling quick adaptation. Integrating these principles into metacognitive learning, reasoning, and control provides the levers and connecting points for instilling generalist behaviors in agents.
@@
The proposed metacognitive framework captures the essence of generalist behaviors and may have far-reaching implications across various domains. In this article, we describe the key components of the framework that exemplify generalist characteristics, followed by an experimental evaluation as well as how we envision it enabling flexibility through meta-level knowledge guided interactions.
@@ 2 Background
In this section, we briefly describe closely related work and the general setting of meta-reasoning, graph neural networks, and policy gradient methods.
Related work
@@
Compositionality and abstractions are central to generalization learning, with abstraction requiring a mechanism for grounding representations in actual objects. In previous work, Veerapaneni et al. (2020) presented an object-centric model where object representations were constructed from raw observations and simpler local functions with uniformity, which were then used for prediction and planning. Similarly, Vijay et al. (2019) used a symbolic object-centric representation to study the impact of adding prior knowledge through relations in an object graph. These studies focus primarily on object space, which inhibits the abstraction of objects and relational structures naturally occurring in the real world and used for decision-making. Key questions include what a natural abstraction should be and how it can be used to compose interactions as building blocks leveraging natural relational structures. We address these largely overlooked questions by preserving abstract knowledge and relational higher-order structures in the form of knowledge graphs, where abstractions are grounded using type space. Learning a compositional structure is achieved through learning over knowledge graphs, with neighborhood aggregation until stability is reached. This distinguishes our approach from previous works.
@@
Other methods emphasize different aspects, such as encouraging exploration through intrinsic reward or information gain Ten et al. (2022); Oudeyer et al. (2007); Pathak et al. (2017), motivation Moerland et al. (2023); Zhu et al. (2023), developing internal goals for empowerment Klyubin et al. (2005), using a curriculum to progress from easier to more difficult tasks Soviany et al. (2022); Campero et al. (2020), or inducing a learning hierarchy based on teacher signals Portelas et al. (2020) Pateria et al. (2021); Hutsebaut-Buysse et al. (2022) or rewards Furelos-Blanco et al. (2023). These methods focus on the execution space, influencing reward signals for learning policies, but largely ignore the higher-order structures and relations between objects in the environment. Our work distinguishes itself by incorporating natural relational structures through knowledge graphs and leverages meta-level knowledge structures to guide interactions that govern low-level execution.
Meta-reasoning
@@
Meta-reasoning is often defined as the set of processes that monitor our thought processes and regulate resource allocation Ackerman and Thompson (2017). Its basic structure Nelson (1990) consists of two levels: meta and object. The meta level’s functions include control and monitoring. The object’s state changes as control passes from the meta to the object level. The monitoring process allows the meta controller to track the status of object-level processes and actions, as well as the consequences of those actions. However, current meta-cognitive reasoning, primarily as a resource manager, is insufficient for exhibiting generalist behaviors, necessitating generalist meta-cognition approaches such as the one proposed here.
Graph neural networks
@@
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is defined by a set of nodes $\mathcal{V}$ with node attributes $\mathtt{X}_{v}$ for $v\in\mathcal{V}$ and a set of edges $\mathcal{E}$ between nodes. Edge from node $u\in\mathcal{V}$ to node $v\in\mathcal{V}$ is given by $(u,v)\in\mathcal{E}$ with attributes $\mathtt{e}_{u,v}$. One typical task is to predict the label or value of the entire graph. Graph neural networks (GNN) (Gilmer et al. (2017); Hamilton et al. (2017) and others) use node connectivity as well as node and edge features to learn a permutation-invariant embedding vector for each node $h_{v},v\in\mathcal{V}$ and the entire graph $h_{\mathcal{G}}$. Typically, the representation is learned through message passing or neighborhood feature aggregation, in which node embedding vectors are updated in each layer based on message aggregation from neighbors:
@@
where $h_{v}^{(k)}$ is the updated representation of node $v$ embedding in layer $k$, $\mathcal{N}(v)$ is the set of neighboring nodes of node $v$, $h_{u}^{(k-1)}$ is the representation of neighboring node $u$ in layer $k-1$, $\bigoplus$ is a differentiable, permutation invariant aggregation function, $\mathcal{F}^{\prime}$ and $\mathcal{F}^{\prime\prime}$ are differentiable functions usually neural networks. Node features are transformed prior to aggregation, and the methods used to transform and aggregate features differ between graph network variants. Motivated by the attention mechanism Vaswani et al. (2017) in the transformer model, Veličković et al. (2017) employs self-attention over node features and aggregating over neighborhood nodes in graph attention networks (GAT). We use the GATv2 operator Brody et al. (2021), which improves static attention problems in graph attention layers.
Policy gradient methods
@@
We consider a standard reinforcement learning setting Sutton and Barto (2018) in which an agent interacts with the environment over a number of steps. A Markov decision process (MDP) is defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ in which $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite discrete action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the transition model, $r:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function and $\gamma\in(0,1)$ is the discount factor. At each step t, the agent takes an action $a_{t}\in\mathcal{A}$ in state $s_{t}\in\mathcal{S}$ according to its policy $\pi$, which is a mapping of states to actions. As a result, the agent reaches state $s_{t+1}\in\mathcal{S}$, receives a scalar reward $r_{t}\in\mathbb{R}$, and accumulates returns; $R_{t}=\sum_{k=0}^{\inf}\gamma^{k}r_{t+k+1}$ represents the accumulated return from step t. The agent’s goal is to maximize the expected return for each state. The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is defined as $Q^{\pi}(s,a)=\mathbb{E}[R_{t}|s_{t}=s,a]$, which is the expected return from selecting action $a$ in state $s_{t}$ following policy $\pi$. The value of state $s$ under policy $\pi$ is defined as $V_{\pi}(s)=\mathbb{E}[R_{t}|s_{t}=s]$, which is the expected return from the state $s$. Policy gradient methods, like REINFORCE Williams (1992) , directly learn a parameterized policy $\pi(s;\theta)$ and update its parameters $\theta$ using approximate gradient ascent of a performance measure, $\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi}[G_{t}\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})]$. To reduce the variance in the expected return estimate, a baseline is subtracted from it; typically, a learned value function is used as the baseline. The resulting quantity $A(s,a)=Q(s,a)-V(s)$ can be viewed as an estimate of the advantage of action $a$ in state $s$. This method is called advantage actor critic (A2C) Sutton and Barto (2018); Degris et al. (2012), and the agent acting according to the policy $\pi$ is the actor, while the value baseline is the critic. Several improved methods have been developed to address sample efficiency, stability, and training difficulty. Proximal policy optimization (PPO) Schulman et al. (2017) is an on-policy policy optimization technique that addresses concerns in actor critic methods using clipped objectives. We adapt PPO for learning policies.
@@ 3 Model
Our proposal for metacognitive generalization centers around the ideas of natural decomposition of tasks that an agent can accomplish through interactions with objects. The meta-policy recommends localized agent-object interactions, which the agent executes following interaction policies to construct a compositional solution. Here, we discuss the main ideas of the framework (see Figure 2) and argue that it exhibits task generalization characteristics.
@@
Figure 2: The meta-cognitive generalization framework consists of three levels: knowledge ($\mathcal{K}$), interaction ($\mathcal{I}$), and execution ($\mathcal{X}$), which are based on abstraction and computation for providing generalization capabilities. The knowledge level contains type space knowledge graphs, which represent higher order abstractions. The instance space contains object instance graphs.The interaction level serves as an intermediary, making recommendations for interactions. The execution level is the lowest level at which the agent acts based on the recommended interaction subgoals. Computations proceed through recognition, recommendation, execution, and interactions to collect experiences for updates to interaction and meta policies.
Knowledge
The environment $\Xi$ is partially observable and consists of objects $e=\{e_{1},e_{2},...,e_{n}\}$ and an agent $A$. It is assumed that the agent has a recognition system that can access its partial view $X$ and associate objects with their corresponding types, such that the type of an object $e_{i}\in e$ is in entity types $E=\{E_{1},E_{2},...,E_{m}\}$. The agent can perform low-level actions $u=\{u_{1},u_{2},...\}$ within the environment. When making decisions, the agent organizes its perceived knowledge into knowledge graphs: an instance-level graph and a type-level graph. An instance-level graph is an ordered pair $G_{I}=(e,\Lambda)$ of observed objects $e$ and their relations $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$. A type-level graph is an ordered pair $G_{K}=(E,\Lambda)$ of entity types $E$ and their type level or abstract relations $\Lambda$. Relations, whether between objects or entity types, retain the same meaning, such as adjacency indicating that two objects are close to one another. The function $\Phi:X\to G_{I}$ maps the agent’s partially observed view to the instance graph by capturing observed objects and deriving their relationships. The function $\Psi:G_{I}\to G_{K}$ maps instance space graphs to type space graphs, a direct translation since the agent’s recognition system contains types corresponding to the objects. If the agent encounters a novel object without a corresponding type, it must create a new type in its type space or assign the object to a related type. In the framework, $\Psi$ is learnable and represents a belief in mapping objects to types. However, this article treats the mapping as fixed, from object nodes in the instance graph $G_{I}$ to type nodes in the type graph $G_{K}$. As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view, resulting in new instance and type space graphs.
The agent’s type space level is also referred to as the meta-level. At the meta-level, the agent uses a meta-policy that accepts a type space knowledge graph and returns meta-level actions or interactions for object interaction. Interactions are represented as $\lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{p}\}$. Similarly, for each interaction type, the agent uses an interaction policy to guide low-level actions $u$ within the environment. Therefore, the agent’s interaction level $\mathcal{I}=(A,G_{I},\lambda)$ comprises of its instance space knowledge graph and available interactions, while the meta-level knowledge $\mathcal{K}=(A,G_{K},\lambda)$ comprises of its type space knowledge graph and interactions.
Interaction
@@
After the agent constructs the knowledge graphs, the meta-level recommender $\Upsilon$ recommends a candidate object $e_{i}$ and interaction type $\lambda_{p}$ for an interaction based on the agent’s knowledge graphs and meta-level policy $\pi$. The recommender iterates through all visible objects in the agent’s instance graph $G_{I}$. For each object, it adds an activation relation between the agent and the object in both the instance space graph $G_{I}$ and the type space graph $G_{K}$, where ‘activating’ refers to observing or focusing on an object for potential interaction. For $G_{K}$, the agent’s meta-policy returns the state value and interaction probability. The state value represents the value of activating that specific object. Then, an object $e_{i}$ for interaction is proposed through sampling based on the values of activation of the visible objects, and interaction $\lambda_{p}$ is sampled from the interaction probability of the proposed object. Therefore, at each step, the meta-level of the agent activates a candidate object $e_{i}$ for interaction $\lambda_{p}$, setting the agent’s interaction goal $g^{i}_{p}=(e_{i},\lambda_{p})$. Acting on the recommended interaction goal, the agent begins interacting with the environment. After each round of interaction, the interaction status is passed to the meta-level, which then proposes the next candidate object and interaction. This process continues until the task is completed or the environment is terminated.
eXecution
Each agent has a single meta-level policy, whereas interaction policies are a collection of policies corresponding to different interactions. By following the interaction goals $g^{i}_{p}$ as directed by the meta-level, the agent interacts with objects. During an interaction $\lambda_{p}$, the agent executes a series of low-level actions $u$ on execution level states $s$ based on the relevant interaction policy $\mu_{p}$. For example, if there is a door, the agent can activate it and interact with it, like ‘open’ the door, even if this requires multiple low-level actions such as moving and toggling. As the agent interacts with objects, it collects trajectories of experiences at both the meta and interaction levels. The meta-level trajectory is used to learn the meta-level policy, while the low-level interaction experiences are used to learn the collection of interaction policies. The learning approach for the meta-policy and interaction policies is detailed below.
Meta-policy learning
@@
Meta-level representations of type space graphs with activation relations are learned using a message-passing graph neural network. Nodes have abstract level attributes such as type encoding, while edges have edge type encoding for different types of relationships. We use GAT convolutional layers Brody et al. (2021); Fey and Lenssen (2019) to learn graph representations given as,
@@
Interaction policy learning
Interaction policies are a repository of policies for different interactions which can be recommended by the meta-level, $\mu_{\phi}=\{\mu_{\phi_{1}},\mu_{\phi_{2}},...,\mu_{\phi_{p}}\}$ and each policy is parameterized separately. When an interaction goal with an object $e_{i}$ and interaction $\lambda_{p}$ is recommended, $g^{i}_{p}=(e_{i},\lambda_{p})$, the agent attempts to achieve it by executing a series of low-level actions from $u$. The agent observes a low-level state $s$ from a convolutional neural network embedding, samples an action $u\sim\mu_{\phi_{p}}(.|s)$ using the respective interaction policy, acts in the environment, and collects the trajectory of experiences $\mathcal{D}_{\tau_{p}}$. The experiences collected are used to learn interaction policies. An estimate of advantages is computed as $A_{p}=r_{p}+\gamma_{p}V_{\mu_{\phi_{p}}}(s^{\prime})-V_{\mu_{\phi_{p}}}(s)$.